### v1

@@ 3 Parabolic dynamics
Combining the Wolff-Denjoy Theorem 1.1 with the Julia-Wolff-Carathéodory theorem for the upper half-plane (see, e.g., [Ab2022, Corollary 2.3.4]) we obtain the following statement:
Theorem 3.1.
Let $F\in\operatorname{\textnormal{Hol}}(\mathbb{H}^{+},\mathbb{H}^{+})$ be a holomorphic self-map without fixed points. Then there exists a point $\tau_{F}\in\partial\mathbb{H}^{+}$ such that $F^{n}\to\tau_{F}$ uniformly on compact subsets. Furthermore, if $\tau_{F}=\infty$ then there exists $F^{\prime}(\tau_{F})\in[1,+\infty)$ such that for any sequence $\{w_{n}\}$ converging non-tangentially to $\infty$ we have
$\lim_{n\to\infty}\frac{F(w_{n})}{w_{n}}=\lim_{n\to\infty}F^{\prime}(w_{n})=F^{\prime}(\infty)\;.$ (3.1)
Here, a sequence $\{w_{n}\}\subset\mathbb{H}^{+}$ converges non-tangentially to $\infty$ if and only if there exists $\varepsilon>0$ such that $\mathop{\textnormal{Im}}w_{n}\geqslant\varepsilon|w_{n}|$ for all $n\in\mathbb{N}$ or, equivalently, if and only if there exists $\delta>0$ such that $\arg w_{n}\in[\delta,\pi-\delta]$ for all $n\in\mathbb{N}$; see, e.g., [Ab2022, Proposition 2.2.7].
@@
The first useful result that we want to prove is the following lemma, which is a generalization (with a proof not depending on models) of [Ab2022, Corollary 4.6.10].
Lemma 3.2.
@@
$\tanh\omega_{\mathbb{H}^{+}}(w_{n},w_{n+1})=\left|\frac{w_{n+1}-w_{n}}{w_{n+1}-\overline{w_{n}}}\right|=\left|\frac{p(w_{n})}{2i\mathop{\textnormal{Im}}w_{n}+p(w_{n})}\right|\leqslant\frac{|p(w_{n})/w_{n}|}{2\frac{\mathop{\textnormal{Im}}w_{n}}{|w_{n}|}-\left|\frac{p(w_{n})}{w_{n}}\right|}\;.$ (3.2)
Since $\{w_{n_{k}}\}$ converges non-tangentially to $\infty$, we can find $\varepsilon>0$ such that $\mathop{\textnormal{Im}}w_{n_{k}}\geqslant\varepsilon|w_{n_{k}}|$ for all $k\in\mathbb{N}$. So (3.2) yields
$\tanh\omega_{\mathbb{H}^{+}}(w_{n_{k}},w_{n_{k}+1})\leqslant\frac{|p(w_{n_{k}})/w_{n_{k}}|}{2\varepsilon-\left|\frac{p(w_{n_{k}})}{w_{n_{k}}}\right|}\;.$ (3.3)
Moreover, (3.1) yields $F(w_{n_{k}})/w_{n_{k}}\to 1$; hence $p(w_{n_{k}})/w_{n_{k}}\to 0$ and thus $\omega_{\mathbb{H}^{+}}(w_{n_{k}},w_{n_{k}+1})\to 0$, by (3.3). But we have already remarked that $\{\omega_{\mathbb{H}^{+}}(w_{n},w_{n+1})\}$ is a non-increasing sequence converging to $s^{F}(w_{0})$; therefore we must have $s^{F}(w_{0})=0$ and, thus, $s^{F}\equiv 0$ by Theorem 2.3. ∎
We can now prove the following statement, that contains the answer to the question posed by Valiron [Val1954, p. 148] mentioned in the introduction in the unit disk setting.
Theorem 3.3.
@@
Put $w_{n}=F^{n}(w_{0})$. Assume, by contradiction, that $\{w_{n}^{-1}F^{n}\}$ does not converge to 1. Then Valiron (see [Val1954, pp. 146-148] and Remark 3.5 below) found a subsequence $\{n_{k}\}$ such that $\{w_{n_{k}}\}$ converges non-tangentially to $\infty$ and such that $\{|w_{n_{k}}|^{-1}F^{n_{k}}\}$ converges to a non-constant map $\psi\in\operatorname{\textnormal{Hol}}(\mathbb{H}^{+},\mathbb{H}^{+})$ satisfying $\psi\circ F=\psi$.
@@
$\geqslant\omega_{\mathbb{H}^{+}}\bigl{(}\psi\bigl{(}F^{n}(z)\bigr{)},\psi\bigl{(}F^{n}(w)\bigr{)}\bigr{)}=\omega_{\mathbb{H}^{+}}\bigl{(}\psi(z),\psi(w)\bigr{)}>0\;.$
@@
Remark 3.4.
The argument used in the proof of the previous theorem is a particular case of [AbSh2025, Corollary 4.6].
Remark 3.5.
For the sake of completeness, we describe here how Valiron in [Val1954, pp. 146-148] produced the subsequence $\{n_{k}\}$ and the non-constant map $\psi$ used in the proof of Theorem 3.3.
@@
In other words, $\{F^{n_{k}}(w)\}$ converges to $\infty$ non-tangentially for all $w\in\mathbb{H}^{+}$. By (3.1), we then have
@@
Corollary 3.6.
Let $f\in\operatorname{\textnormal{Hol}}(\mathbb{D},\mathbb{H}^{+})$ be parabolic with Wolff point $\tau_{f}$. Fix $z_{0}\in\mathbb{D}$. Then
@@
Up to conjugating by a rotation, without loss of generality we can assume $\tau_{f}=1$. Let $F=\Psi\circ f\circ\Psi^{-1}$, where $\Psi$ is the Cayley transform. Then $F$ is a parabolic self-map of $\mathbb{H}^{+}$ with Wolff point at $\infty$. Setting $w_{0}=\Psi(z_{0})$, we have $f^{n}(z_{0})=\Psi^{-1}\bigl{(}F^{n}(w_{0})\bigr{)}$ and
$\frac{f^{n}(z)-1}{f^{n}(z_{0})-1}=\frac{F^{n}(w_{0})+i}{F^{n}\bigl{(}\Psi(z)\bigr{)}+i}=\frac{F^{n}(w_{0})}{F^{n}\bigl{(}\Psi(z)\bigr{)}}\frac{1+i/F^{n}(w_{0})}{1+i/F^{n}\bigl{(}\Psi(z)\bigr{)}}\;$;
so the assertion follows from Theorem 3.3. ∎
@@
Take $z\in\mathbb{D}$. By Corollary 3.6 we know that $\frac{f^{n}(z)-\tau_{f}}{f^{n}(z_{0})-\tau_{f}}\to 1$; therefore,
@@
$\frac{1-z}{|1-z|}=\frac{r^{-1}+ie^{-i\theta}}{\sqrt{r^{-2}+1}}\;.$
In particular, $w$ tends to $\infty$ in such a way that $\arg w$ converges to $\phi\in[0,\pi]$ if and only if $z\to 1$ in such a way that $\frac{z-1}{|z-1|}\to-ie^{-i\phi}$. Then the assertion follows from Corollary 3.7 applied to $f=\Psi^{-1}\circ F\circ\Psi$, which is parabolic with Wolff point $1$. ∎
This result does not imply that the orbits of a parabolic map always converge to the Wolff point along a given slope. Indeed, there are examples of parabolic maps with zero hyperbolic step whose orbits do not converge to the Wolff point tangentially to some direction; see, e.g., [Braccietalbook, Example 17.5.4].
On the other hand, if $f$ is parabolic with positive hyperbolic step, by Lemma 3.2 no orbit can have a subsequence converging non-tangentially to the Wolff point. However, this does not immediately implies that all orbits converge to the Wolff point with the same slope, because in $\tau_{f}$ there are two tangential rays and, in principle, an orbit might jump from one ray to the other. Luckily, this does not happen. This result has already been proved by Pommerenke [Pomm1979]; but using the ideas presented so far we can give an easier proof.
Theorem 3.9.
@@
$\lim_{n\to+\infty}\left|F^{n}(w)-\pi/2\right|=\pi/2$ (3.4)
for all $w\in\mathbb{H}^{+}$. If this would not be true, there would exist $w_{0}\in\mathbb{H}^{+}$ and a subsequence $\{F^{n_{k}}\}$ such that $\sup_{k}|\arg F^{n_{k}}(w_{0})-\pi/2|<\pi/2$. But then $F^{n_{k}}(w_{0})\to\infty$ non-tangentially and then, by Lemma 3.2, we would have $s^{F}\equiv 0$, impossible.
Now, fix $w_{0}\in\mathbb{H}^{+}$. We claim that either $\arg F^{n}(w_{0})\to 0$ or $\arg F^{n}(w_{0})\to\pi$. If this were not true, by (3.4) we could find subsequences $\{F^{\mu_{k}}\}$ and $\{F^{\nu_{k}}\}$ such that $\arg F^{\mu_{k}}(w_{0})\to 0$ and $\arg F^{\nu_{k}}(w_{0})\to\pi$; up to taking further subsequences, we can also assume $\mu_{k}<\nu_{k}<\mu_{k+1}$ for all $k\in\mathbb{N}$. We can now define a continuous curve $\sigma\colon[0,+\infty)\to\mathbb{H}^{+}$ as follows:
$\sigma(t)=\begin{cases}(1-t)w_{0}+tF(w_{0})&\hbox{if $0\leqslant t\leqslant 1$};\\ F^{n+1}\bigl{(}\sigma(t-(n+1))\bigr{)}&\hbox{if $n+1\leqslant t\leqslant n+2$ and $n\in\mathbb{N}$}.\end{cases}$
Then $\arg\sigma$ is a continuous function defined on the connected set $0,+\infty)$ and whose image contains both values converging to 0 and values converging to $\pi$; therefore we can find a sequence $\{t_{k}\}\subset[0,1]$ and a subsequence $\{F^{n_{k}}\}$ such that $\arg F^{n_{k}}\bigl{(}\sigma(t_{k})\bigr{)}=\pi/2$ for all $k\in\mathbb{N}$. Up to taking further subsequences, we can assume that $t_{k}\to\tau\in[0,1]$ and that $\arg F^{n_{k}}$ converges uniformly on $\sigma([0,1])$. But then we get $\lim\limits_{k\to+\infty}\arg F^{n_{k}}(\tau)=\pi/2$, against (3.4).
So for each $w_{0}\in\mathbb{H}^{+}$ we have either $\arg F^{n}(w_{0})\to 0$ or $\arg F^{n}(w_{0})\to\pi$. But, by Corollary 3.8, all points must converge to $\infty$ with the same argument and we are done. ∎
@@
Up to conjugating by a rotation, we can without loss of generality assume that $\tau_{f}=1$. The assertion then follows by Theorem 3.9 arguing as in the proof of Corollary 3.8, reversing the roles of $f$ and $F$. ∎

### v2

@@ 3 Parabolic dynamics
Combining the Wolff-Denjoy Theorem 1 with the Julia-Wolff-Carathéodory theorem for the upper half-plane (see, e.g., [4, Corollary 2.3.4]) we obtain the following statement:
Theorem 5.
Let $F\in\operatorname{\textnormal{Hol}}(\mathbb{H}^{+},\mathbb{H}^{+})$ be a holomorphic self-map without fixed points. Then there exists a point $\tau_{F}\in\partial\mathbb{H}^{+}$ such that $F^{n}\to\tau_{F}$ uniformly on compact subsets. Furthermore, if $\tau_{F}=\infty$ then there exists a real number $F^{\prime}(\infty)\geq 1$ such that for any sequence $\{w_{n}\}$ converging non-tangentially to $\infty$ we have
$\lim_{n\to\infty}\frac{F(w_{n})}{w_{n}}=\lim_{n\to\infty}F^{\prime}(w_{n})=F^{\prime}(\infty)\;.$ (5)
Here, a sequence $\{w_{n}\}\subset\mathbb{H}^{+}$ converges non-tangentially to $\infty$ if and only if there exists $\varepsilon>0$ such that $\mathop{\textnormal{Im}}w_{n}\geq\varepsilon|w_{n}|$ for all $n\in\mathbb{N}$ or, equivalently, if and only if there exists $\delta>0$ such that $\arg w_{n}\in[\delta,\pi-\delta]$ for all $n\in\mathbb{N}$; see, e.g., [4, Proposition 2.2.7].
@@
The first useful result that we want to prove is the following lemma, which is a generalization (and with a proof not depending on models) of [4, Corollary 4.6.10].
Lemma 6.
@@
$\tanh\omega_{\mathbb{H}^{+}}(w_{n},w_{n+1})=\left|\frac{w_{n+1}-w_{n}}{w_{n+1}-\overline{w_{n}}}\right|=\left|\frac{p(w_{n})}{2i\mathop{\textnormal{Im}}w_{n}+p(w_{n})}\right|\leq\frac{|p(w_{n})/w_{n}|}{2\frac{\mathop{\textnormal{Im}}w_{n}}{|w_{n}|}-\left|\frac{p(w_{n})}{w_{n}}\right|}\;.$ (6)
Since $\{w_{n_{k}}\}$ converges non-tangentially to $\infty$, we can find $\varepsilon>0$ such that $\mathop{\textnormal{Im}}w_{n_{k}}\geq\varepsilon|w_{n_{k}}|$ for all $k\in\mathbb{N}$. So (6) yields
$\tanh\omega_{\mathbb{H}^{+}}(w_{n_{k}},w_{n_{k}+1})\leq\frac{|p(w_{n_{k}})/w_{n_{k}}|}{2\varepsilon-\left|\frac{p(w_{n_{k}})}{w_{n_{k}}}\right|}\;.$ (7)
Moreover, (5) yields $F(w_{n_{k}})/w_{n_{k}}\to 1$; hence $p(w_{n_{k}})/w_{n_{k}}\to 0$ and thus $\omega_{\mathbb{H}^{+}}(w_{n_{k}},w_{n_{k}+1})\to 0$, by (7). But we have already remarked that $\{\omega_{\mathbb{H}^{+}}(w_{n},w_{n+1})\}$ is a non-increasing sequence converging to $s^{F}(w_{0})$; therefore we must have $s^{F}(w_{0})=0$ and, thus, $s^{F}\equiv 0$ by Theorem 4. ∎
We can now prove the following statement, that contains the answer to the question posed by Valiron [7, p. 148] mentioned in the introduction in the unit disk setting.
Theorem 7.
@@
Put $w_{n}=F^{n}(w_{0})$. Assume, by contradiction, that $\{w_{n}^{-1}F^{n}\}$ does not converge to 1. Then Valiron (see [7, pp. 146-148] and Remark 2 below) found a subsequence $\{n_{k}\}$ such that $\{w_{n_{k}}\}$ converges non-tangentially to $\infty$ and such that $\{|w_{n_{k}}|^{-1}F^{n_{k}}\}$ converges to a non-constant map $\psi\in\operatorname{\textnormal{Hol}}(\mathbb{H}^{+},\mathbb{H}^{+})$ satisfying $\psi\circ F=\psi$.
@@
$\geq\omega_{\mathbb{H}^{+}}\bigl(\psi\bigl(F^{n}(z)\bigr),\psi\bigl(F^{n}(w)\bigr)\bigr)=\omega_{\mathbb{H}^{+}}\bigl(\psi(z),\psi(w)\bigr)>0\;.$
@@
Remark 1.
The argument used in the proof of the previous theorem is a particular case of [5, Corollary 4.6].
Remark 2.
For the sake of completeness, we describe here how Valiron in [7, pp. 146-148] produced the subsequence $\{n_{k}\}$ and the non-constant map $\psi$ used in the proof of Theorem 7.
@@
In other words, $\{F^{n_{k}}(w)\}$ converges to $\infty$ non-tangentially for all $w\in\mathbb{H}^{+}$. By (5), we then have
@@
Corollary 8.
Let $f\in\operatorname{\textnormal{Hol}}(\mathbb{D},\mathbb{D})$ be parabolic with Wolff point $\tau_{f}$. Fix $z_{0}\in\mathbb{D}$. Then
@@
Up to conjugating by a rotation, without loss of generality we can assume $\tau_{f}=1$. Let $F=\Psi\circ f\circ\Psi^{-1}$, where $\Psi$ is the Cayley transform. Then $F$ is a parabolic self-map of $\mathbb{H}^{+}$ with Wolff point at $\infty$. Recalling that $\Psi^{-1}(w)=\frac{w-i}{w+i}$, for any $z\in\mathbb{D}$ we find
$f^{n}(z)-1=\frac{F^{n}\bigl(\Psi(z)\bigr)-i}{F^{n}\bigl(\Psi(z)\bigr)+i}-1=\frac{-2i}{F^{n}\bigl(\Psi(z)\bigr)+i}$
and hence
$\frac{f^{n}(z)-1}{f^{n}(z_{0})-1}=\frac{F^{n}(w_{0})+i}{F^{n}\bigl(\Psi(z)\bigr)+i}=\frac{F^{n}(w_{0})}{F^{n}\bigl(\Psi(z)\bigr)}\frac{1+i/F^{n}(w_{0})}{1+i/F^{n}\bigl(\Psi(z)\bigr)}\;,$
where $w_{0}=\Psi(z_{0})$. So the assertion follows from Theorem 7. ∎
@@
Take $z\in\mathbb{D}$. By Corollary 8 we know that $\frac{f^{n}(z)-\tau_{f}}{f^{n}(z_{0})-\tau_{f}}\to 1$; therefore,
@@
$\frac{1-z}{|1-z|}=\frac{r^{-1}+ie^{-i\theta}}{|r^{-1}+ie^{-i\theta}|}\;.$
In particular, $w$ tends to $\infty$ in such a way that $\arg w$ converges to $\phi\in[0,\pi]$ if and only if $z\to 1$ in such a way that $\frac{z-1}{|z-1|}\to-ie^{-i\phi}$. Then the assertion follows from Corollary 9 applied to $f=\Psi^{-1}\circ F\circ\Psi$, which is parabolic with Wolff point $1$. ∎
This result does not imply that the orbits of a parabolic map always converge to the Wolff point along a given slope. Indeed, there are examples of parabolic maps with zero hyperbolic step whose orbits do not converge to the Wolff point tangentially to some direction; see, e.g., [13, Example 17.5.4].
On the other hand, if $f$ is parabolic with positive hyperbolic step, by Lemma 6 no orbit can have a subsequence converging non-tangentially to the Wolff point. However, this does not immediately imply that all orbits converge to the Wolff point with the same slope, because in $\tau_{f}$ there are two tangential rays and, in principle, an orbit might jump from one ray to the other. Luckily, this does not happen. This result has already been proved by Pommerenke [8]; but using the ideas presented so far we can give an easier proof.
Theorem 11.
@@
$\lim_{n\to+\infty}\left|F^{n}(w)-\pi/2\right|=\pi/2$ (8)
for all $w\in\mathbb{H}^{+}$. If this would not be true, there would exist $w_{0}\in\mathbb{H}^{+}$ and a subsequence $\{F^{n_{k}}\}$ such that $\sup_{k}|\arg F^{n_{k}}(w_{0})-\pi/2|<\pi/2$. But then $F^{n_{k}}(w_{0})\to\infty$ non-tangentially and then, by Lemma 6, we would have $s^{F}\equiv 0$, impossible.
Now, fix $w_{0}\in\mathbb{H}^{+}$. We claim that either $\arg F^{n}(w_{0})\to 0$ or $\arg F^{n}(w_{0})\to\pi$. If this were not true, by (8) we could find subsequences $\{F^{\mu_{k}}\}$ and $\{F^{\nu_{k}}\}$ such that $\arg F^{\mu_{k}}(w_{0})\to 0$ and $\arg F^{\nu_{k}}(w_{0})\to\pi$; up to taking further subsequences, we can also assume $\mu_{k}<\nu_{k}<\mu_{k+1}$ for all $k\in\mathbb{N}$. We can now define a continuous curve $\sigma\colon[0,+\infty)\to\mathbb{H}^{+}$ as follows:
$\sigma(t)=\begin{cases}(1-t)w_{0}+tF(w_{0})&\hbox{if $0\leq t\leq 1$};\\ F^{n+1}\bigl(\sigma(t-(n+1))\bigr)&\hbox{if $n+1\leq t\leq n+2$ and $n\in\mathbb{N}$}.\end{cases}$
Then $\arg\sigma$ is a continuous function defined on the connected set $0,+\infty)$ and whose image contains both values converging to 0 and values converging to $\pi$; therefore we can find a sequence $\{t_{k}\}\subset[0,1]$ and a subsequence $\{F^{n_{k}}\}$ such that $\arg F^{n_{k}}\bigl(\sigma(t_{k})\bigr)=\pi/2$ for all $k\in\mathbb{N}$. Up to taking further subsequences, we can assume that $t_{k}\to\tau\in[0,1]$ and that $\arg F^{n_{k}}$ converges uniformly on $\sigma([0,1])$. But then we get $\lim\limits_{k\to+\infty}\arg F^{n_{k}}(\tau)=\pi/2$, against (8).
So for each $w_{0}\in\mathbb{H}^{+}$ we have either $\arg F^{n}(w_{0})\to 0$ or $\arg F^{n}(w_{0})\to\pi$. But, by Corollary 10, all points must converge to $\infty$ with the same argument and we are done. ∎
@@
Up to conjugating by a rotation, we can without loss of generality assume that $\tau_{f}=1$. The assertion then follows by Theorem 11 arguing as in the proof of Corollary 10, reversing the roles of $f$ and $F$. ∎